OpenStudy (anonymous):

The temperature was 65 degrees at daybreak. Then it dropped two degrees per hour until dusk. If x is the number of hours since daybreak, this equation represents the temperature during the day: y = -2x + 65 OR y = 65 - 2x. Use the equation to predict the temperature six hours later. Answer 53 63 61 59

OpenStudy (anonymous):

plug 6 in x, gives 53

OpenStudy (anonymous):

yep^^

OpenStudy (anonymous):

replace X with 6 which would be y=65-2(6) then y=65-12which will give u 52

OpenStudy (anonymous):

that 52 surpose to be 53